# LlamaRestTest: Effective REST API Testing with Small Language Models

### v1

@@ 4.3. RQ2: Impact of Fine-Tuning and Quantization in REST API Testing
Table 5. Average Number of Operations Processed Across 10 Runs for Closed Source Services Using LlamaRestTest with Llama3-8B (Vanilla), Fine-Tuning (FT), and Quantization (Q)
Service FT with Q(8-Bit) FT with Q(4-Bit) FT with Q(2-Bit) FT Vanilla
@@
Ohsome 24.5 21.4 13.2 12.1 11.2
Sum 37.5 34.4 25.4 24.5 23.0
Table 6. Average Number of Operations Processed for Closed Source Services Across Various REST API Testing Tools (LlamaRestTest with FT with Q(8-Bit) Configuration)
@@
Ohsome 0.0 0.0 0.0 0.0 24.5
Sum 9.9 10.6 11.6 11.2 37.5
Figure 3. Code Coverage Achieved Across 10 Runs for Open Source Services Using LlamaRestTest with Llama3-8B (Vanilla), Fine-Tuning (FT), and Quantization (Q)
@@
To evaluate the impact of fine-tuning and various quantization levels on REST API testing, we assessed LlamaRestTest under five configurations: the vanilla Llama3-8B model, a fine-tuned version, and quantized versions at 8-bit, 4-bit, and 2-bit. The evaluation was conducted across 12 services, including nine open-source and three closed-source services, as detailed in Table 2. The key metrics assessed were method coverage, branch coverage, line coverage, and operation coverage, which are widely recognized as the most popular metrics for measuring the effectiveness of REST API testing tools (Golmohammadi et al., 2023; Kim et al., 2022).
@@
Figure 3 presents a comparison of method, branch, and line coverage across multiple services for the five LlamaRestTest configurations: the fine-tuned version (F), and fine-tuned versions with 8-bit, 4-bit, and 2-bit quantization, alongside the vanilla model. Each panel represents one of the coverage metrics—method, branch, or line—across the various services. The fine-tuned and quantized versions generally exhibit improved coverage compared to the vanilla model, with significant gains observed for services such as Language-Tool and YouTube-Mock. The average coverage for each configuration is also displayed.
@@
The result shows that LlamaRestTest with 8-bit quantization exhibited a 13.4% improvement in method coverage compared to the Vanilla model. The 4-bit quantized version resulted in a 9.8% improvement, while the 2-bit quantized version delivered a 7.5% increase. For branch coverage, the 8-bit quantized version of LlamaRestTest demonstrated a 21.9% improvement over Vanilla. The 4-bit quantized model showed a 20.3% gain, while the 2-bit quantized version achieved a 10.8% improvement. For line coverage, the improvements were more moderate but still significant. On average, the 8-bit quantized version of LlamaRestTest showed an 11.2% improvement over the Vanilla model. The 4-bit quantized model improved line coverage by 10.3%, while the 2-bit quantized model exhibited an 8.5% increase.
@@
Table 6 shows the number of successfully processed operations for closed-source services with the five LlamaRestTest configurations.hows the absolute number of successfully processed operations across closed-source services for the five LlamaRestTest configurations. When comparing the results, LlamaRestTest with 8-bit quantization processed 37.5 operations, representing a 63% increase compared to the Vanilla model, which processed 23 operations. The 4-bit quantized version processed 34.4 operations, a 49.6% improvement over Vanilla, while the 2-bit version processed 25.4 operations, showing a 10.4% relative increase.
The results demonstrate that both effectiveness and efficiency are critical factors in REST API testing. LlamaRestTest with Fine-Tuning with 8-bit quantization achieved the highest code and operation coverage. For example, in RQ1’s IPD rule extraction, the fine-tuned model required 48.9 seconds per IPD on average, while the 8-bit quantized version reduced that to 36.9 seconds. The 4-bit and 2-bit quantized models further reduced the time to 26.2 and 26.1 seconds, respectively. The smaller models benefit from faster inference and model loading times, which enhance the effectiveness of REST API testing by refining responses with server messages in real time.
LlamaRestTest with fine-tuning and 8-bit quantization achieved the highest code and operation coverage, demonstrating improvements of 13.4%, 21.9%, 11.2%, and 63% in method, branch, line, and operation coverage, respectively, while maintaining a strong balance between efficiency and effectiveness.
@@ 4.4. RQ3: Comparison with State-of-the-art REST API Testing Tools
Figure 4. Code Coverage Achieved Across 10 Runs for Open Source Services with MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest (Fine-Tuning with 8-Bit Quantized)
Using the same services from Section 4.3, we measured operation coverage and code coverage, including branch, line, and method coverage, to compare the effectiveness of REST API testing with state-of-the-art tools. Figure 4 presents the code coverage achieved by MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest across various open-source REST APIs in our benchmark, showing the percentages of branch, line, and method coverage. In this experiment, the testing tools used enhanced version of specification generated by RESTGPT.
The results reveal that LlamaRestTest consistently outperforms all other tools across method, branch, and line coverage metrics. For method coverage, LlamaRestTest achieved 55.8%, which is a 67.6% relative improvement over MoRest (33.3%), a 68.6% improvement over RESTler (33.1%), a 33.8% improvement over ARAT-RL (41.7%), and a 21.8% improvement over EvoMaster (45.8%). In branch coverage, LlamaRestTest showed the most significant gains, achieving 28.3%. This represents a 122.8% improvement over MoRest (12.7%), a 195.3% improvement over RESTler (9.6%), a 48.2% improvement over ARAT-RL (19.1%), and a 58.9% improvement over EvoMaster (17.8%). For line coverage, LlamaRestTest achieved 55.3%, marking a 59.9% improvement over MoRest (34.6%), a 66.6% improvement over RESTler (33.2%), a 27.4% improvement over ARAT-RL (43.4%), and a 22.1% improvement over EvoMaster (45.3%).
@@
Table 6 presents the operation coverage, showing the rounded average number of successful operations (Op) achieved by each tool for the closed-source services. Similar with the code coverage, LlamaRestTest outperformed the other tools by a wide margin, with a total score of 37.5. This represents a relative improvement of 278.8% over MoRest (9.9), 254.7% over RESTler (10.6), 223.3% over ARAT-RL (11.6), and 234.8% over EvoMaster (11.2). These results demonstrate that leveraging server response feedback with LlamaREST-IPD and LlamaREST-EX significantly improves testing effectiveness compared to current state-of-the-art tools, even when using enhanced specifications generated by RESTGPT.
LlamaRestTest significantly outperforms other API testing tools, achieving 48.2%-195.3% more branch coverage, 22.1%-66.6% more line coverage, 21.8%-68.6% more method coverage, and processing 223.3%-278.8% more operations.
@@ 4.5. RQ4: Fault-Detection Capability
Table 2 presents the fault-detection capability of five REST API testing tools—MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest—measured by the number of faults detected across all services in our benchmark. We used the 500 internal server error counting script provided by the ARAT-RL artifact, which automatically counts the 500 errors in the server log and discards duplicate entries. The sum of faults detected by each tool illustrates their overall effectiveness in identifying bugs across a range of open-source services.
LlamaRestTest demonstrated the highest fault-detection capability, with a total score of 204. This represents a 45.7% improvement over MoRest (140), a 56.9% improvement over RESTler (130), a 27.5% improvement over ARAT-RL (160), and a 56.9% improvement over EvoMaster (130). The results clearly show that LlamaRestTest outperforms all other tools, detecting more faults across diverse services.
The most notable detected errors were in the Ohsome and Spotify cases. For Ohsome, LlamaRestTest was the only tool to detect faults, identifying 36 bugs where all other tools failed. Similarly, LlamaRestTest was the only tool to find faults in Spotify, detecting 8 errors where all others found none. Both cases involve processing complex scenarios. For example, the get /playlists/playlist_id/tracks operation in Spotify’s API requires specific knowledge of how Spotify generates playlist_id. Spotify generates IDs for playlists that are typically 22 characters long, with constraints on the permitted characters and patterns. While many tools fail to generate valid IDs, LlamaREST-EX uses its LLM to accurately retrieve and generate valid Spotify playlist IDs. As a result, LlamaRestTest is able to successfully query the get /playlists/playlist_id/tracks operation, triggering a chain of interactions across the service. For instance, after retrieving the International Standard Recording Code (ISRC) from the playlist’s tracks, the mutation process randomly selects an ISRC to use as the user_id in the subsequent get /users/user_id/playlists operation. This sequence exposes a hidden dependency conflict, causing an Internal Server Error.
LlamaRestTest surpasses other tools in fault-detection by averaging 204 Internal Server Error (500) responses per run, identifying 44 to 74 more bugs than state-of-the-art REST API testing tools and highlighting its advanced capability to detect internal server errors.
@@ 4.6. RQ5: Ablation Study
Table 7. Ablation Study on the Impact of Key Components on Coverage Metrics
@@
1. Without Parameter Description 49.9% (-10.6%) 26.9% (-4.9%) 50.8% (-8.1%)
2. Without Server Response 49.7% (-10.9%) 23.9% (-15.5%) 49.0% (-11.4%)
3. Without LlamaREST-IPD 46.6% (-16.5%) 22.1% (-21.9%) 46.9% (-15.2%)
4. Without LlamaREST-EX 52.1% (-6.6%) 23.8% (-14.7%) 52.3% (-5.4%)
To assess the contribution of each key component in LlamaRestTest, we conducted an ablation study by systematically removing specific features and measuring the resulting impact on method, branch, and line coverage across three runs. Table 7 presents the coverage achieved by LlamaRestTest with and without these components, allowing us to evaluate how much each feature contributes to the overall performance.
The full LlamaRestTest system achieved method coverage of 55.8%, branch coverage of 28.3%, and line coverage of 55.3%. Removing the parameter description feature resulted in a noticeable drop, with method coverage decreasing by 10.6%, branch coverage by 4.9%, and line coverage by 8.1%. This suggests that including parameter descriptions improves test case generation by providing better information to generate valid inputs. Interestingly, excluding the server response feature caused an even more significant drop, particularly in branch coverage, which decreased by 15.5%. Method and line coverage also fell by 10.9% and 11.4%, respectively. While previous studies in this area have focused primarily on enhancing specifications, these results highlight the critical role of server feedback in guiding test case generation and increasing coverage by detecting dynamic behaviors during runtime.
The removal of the LlamaREST-IPD (Inter-Parameter Dependency) module had the most substantial impact on performance, reducing method coverage by 16.5%, branch coverage by 21.9%, and line coverage by 15.2%. This suggests that modeling inter-parameter dependencies is crucial for generating more effective test cases that cover complex interactions between parameters. Lastly, excluding LlamaREST-EX resulted in moderate reductions in coverage, with method, branch, and line coverage decreasing by 6.6%, 14.7%, and 5.4%, respectively. While LlamaREST-EX still contributes meaningfully to overall test effectiveness, its impact appears to be less critical than that of the other components.
The ablation study shows that server response handling and inter-parameter dependency modeling are the most critical components of LlamaRestTest, with their removal causing the largest drops in coverage, while all components contribute to the overall performance.

### v2

@@ 4.3. RQ2: Impact of Fine-Tuning and Quantization on REST API Testing
Table 5. Average number of operations processed in 10 runs for closed-source services using LlamaRestTest configured with Llama3-8B (Base), fine-tuning (FT), and quantization (Q).
Service FT with 8-Bit Q FT with 4-Bit Q FT with 2-Bit Q FT Base
@@
Total 37.5 34.4 25.4 24.5 23.0
Table 6. Average number of operations processed for closed-source services by different testing tools (LlamaRestTest configured with 8-bit quantization).
@@
Total 9.9 10.6 11.6 11.2 37.5
@@
To evaluate the impact of fine-tuning and various quantization levels on REST API testing, we assessed LlamaRestTest under five configurations: the base Llama3-8B model, the fine-tuned version (at full precision), and the quantized versions at 8-bit, 4-bit, and 2-bit quantization. We evaluated the effectiveness of these configurations in terms of code coverage achieved for the open-source services and operations processed for the closed-source services in our benchmark (Table 2).
@@
Figure 3 presents a comparison of method, branch, and line coverage across multiple services for the five LlamaRestTest configurations: the fine-tuned version, the fine-tuned versions with 8-bit, 4-bit, and 2-bit quantization, and the base model. Each panel represents one of the coverage metrics—method, branch, or line—for the nine open-source services in our benchmark. The figure also shows the average coverage across services for each configuration. The fine-tuned and quantized versions generally exhibited higher coverage compared to the base model, with significant gains observed for the Language-Tool and YouTube services in particular.
The results show that fine-tuning alone produced modest improvements over the base model. The fine-tuned model achieved average coverage rates of 50.11% for method coverage (compared to 49.21% for base), 23.55% for branch coverage (compared to 23.24% for base), and 49.91% for line coverage (compared to 48.51% for base). This represents improvements of approximately 0.9, 0.3, and 1.4 percentage points respectively, demonstrating that fine-tuning provides an observable but limited enhancement to testing capability.
@@
On combining fine-tuning with quantization, we observed substantially greater improvements. LlamaRestTest with 8-bit quantization exhibited 6.6 percentage points improvement in method coverage compared to the base model, reaching 55.8% coverage on average. The 4-bit quantized version resulted in a coverage gain of 4.8 percentage points (54.0% method coverage), while the 2-bit quantized version achieved 3.7 more percentage points (52.9% method coverage). The trends are similar for branch and line coverage, with the quantized versions performing better than the fine-tuned model in gains over the base model. In terms of branch coverage, the 8-bit, 4-bit, and 2-bit quantized versions showed increases of 5.1, 4.7, and 2.5 percentage points, respectively, over the base model; likewise, for line coverage, the gains of these quantized versions were 6.8, 5.0, and 4.1 percentage points, respectively.
@@
Table 6 shows the number of successfully processed operations for the closed-source services in our benchmark under the five LlamaRestTest configurations. LlamaRestTest with fine-tuning alone processed 24.5 operations, which is 1.5 operations more than the base model’s 23 operations. More significantly, LlamaRestTest with 8-bit quantization processed 37.5 operations (14.5 operations more than the base model’s 23), while the 4-bit quantized version processed 34.4 operations (11.4 operations more than the base model). Finally, the 2-bit version was also more effective than the base model, processing 25.4 operations (2.4 operations more than the base model).
These results demonstrate that fine-tuning alone provides modest benefits, but the combination of fine-tuning with appropriate quantization delivers significantly better performance. There are two main contributing factors for this phenomenon. First, the faster inference times of quantized models enable more test generation within the same testing budget, processing for instance 22–81% more inference tasks in the same time budget (Dettmers et al., 2022). For example, for IPD rule extraction (RQ1), the fine-tuned model required 48.9 seconds per IPD on average, whereas the 8-bit quantized version reduced the inference time to 36.9 seconds. The 4-bit and 2-bit quantized models further reduced the time to 26.2 and 26.1 seconds, respectively. The smaller models benefit from faster inference, which enhances the effectiveness of REST API testing via more efficient processing of server response messages. Second, quantization introduces a regularization effect that may help the model generalize better with various answers rather than memorizing specific examples (Gholami et al., 2022).
Although quantized models generally performed well, we observed scenarios where the 2-bit quantized model underperformed compared to the non-quantized model. Specifically, for services, such as Genome-Nexus and Spotify, that rely heavily on domain-specific knowledge, the 2-bit quantized model often generated incorrect outputs, leading to lower code coverage. This suggests that the extreme compression in the 2-bit model may result in a loss of precision that affects performance in complex, domain-specific scenarios. However, the 4-bit and 8-bit quantized models consistently outperformed the non-quantized models, even in these challenging cases. This indicates that these quantization levels strike a good balance between model compression and maintaining the necessary precision for effective API testing across various domains.
LlamaRestTest with fine-tuning and 8-bit quantization achieved the highest code and operation coverage, processing 37.5 operations while achieving improvements of 5.5, 5.1, and 6.8 percentage points in method, branch, and line coverage, respectively, over the base model, thus showing a good balance between efficiency and effectiveness. Fine-tuning alone provided modest gains over the base model. While 4-bit and 8-bit quantized models consistently outperformed non-quantized versions, the 2-bit model showed limitations in complex, domain-specific scenarios, highlighting the trade-off between model compression and precision for REST API testing.
@@ 4.4. RQ3: Comparison with State-of-the-art REST API Testing Tools
Figure 4. Code coverage achieved on 10 runs for open-source services with MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest (fine-tuned with 8-bit quantization).
[…]
Code Coverage Achieved Across 10 Runs for Open Source Services with MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest (Fine-Tuning with 8-Bit Quantized)
Next, we compared LlamaRestTest against four state-of-the-art REST API testing tools in terms of code coverage (for the open-source services) and operation coverage (for the closed-source services). For this study, the testing tools used enhanced versions of OpenAPI specifications of the services generated by RESTGPT (Kim et al., 2024b).
Figure 4 presents the code coverage achieved by MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest on the open-source services in our benchmark. The results show that LlamaRestTest consistently outperforms all other tools in method, branch, and line coverage. For method coverage, LlamaRestTest achieved 55.8%, exceeding MoRest (33.3%) by 22.5 percentage points, RESTler (33.1%) by 22.7 percentage points, ARAT-RL (41.7%) by 14.1 percentage points, and EvoMaster (45.8%) by 10 percentage points. In terms of branches, LlamaRestTest reached 28.3% coverage, achieving 15.6 increase over MoRest, 18.7 increase over RESTler, 9.2 increase over ARAT-RL, and 10.5 increase over EvoMaster. Finally, for line coverage, LlamaRestTest achieved 55.3% coverage, again outperforming MoRest (+20.7), RESTler (+22.1), and ARAT-RL (+11.9), and EvoMaster (+10) considerably.
@@
Table 6 presents results on operation coverage, showing the average number of operations processed by each tool for the closed-source services. Like the results on code coverage, LlamaRestTest outperformed the other tools by a wide margin, successfully processing 37.5 operations overall. This represents improvements of 27.6 operations over MoRest (9.9), 26.9 operations over RESTler (10.6), 25.9 operations over ARAT-RL (11.6), and 26.3 operations over EvoMaster (11.2).
LlamaRestTest’s effective performance comes from its ability to analyze server responses, which allows it to detect errors that other tools can miss. For example, in testing Ohsome and Spotify APIs, LlamaRestTest was able to process more operations by utilizing information from server messages. In the case of Ohsome, where the server returned a message indicating that users must specify only one of three parameters, LlamaRestTest used this information to test the operations that the other tools missed. Conversely, LlamaRestTest may be less effective on services that provide minimal or poor server feedback. In such cases, LlamaRestTest could spend time analyzing response messages that do not contain useful information for refining requests, potentially reducing its efficiency.
These results demonstrate that leveraging server response feedback with LlamaREST-IPD and LlamaREST-EX significantly improves testing effectiveness compared to current state-of-the-art tools, even when using enhanced specifications generated by RESTGPT. The ability to interpret and utilize server responses gives LlamaRestTest an edge in most scenarios, particularly those with informative server feedback.
LlamaRestTest significantly outperforms other API testing tools, achieving 10.0–22.7 percentage points higher method coverage, 9.2–18.7 percentage points higher branch coverage, 10.0–22.1 percentage points higher line coverage, and processing 2.5x–3x more operations than competing tools. It performs particularly well in scenarios with informative server feedback, but may be less effective—performing similarly to other tools—on APIs with response messages that lack useful information.
@@ 4.5. RQ4: Fault-Detection Capability
Table 2 presents our results on fault-detection capability of five REST API testing tools—MoRest, RESTler, ARAT-RL, EvoMaster, and LlamaRestTest—measured as the number of 500 internal server errors detected across all services in our benchmark. To measure unique server errors, we used the error-counting script provided in the ARAT-RL artifact (Kim, 2024), which automatically counts the 500 errors in the server log and discards duplicate entries.
LlamaRestTest demonstrated the highest fault-detection ability, discovering 204 faults in 10 runs on the services in our benchmark. This represents 64 more faults detected than MoRest (140), 74 more than RESTler (130), 44 more than ARAT-RL (160), and 74 more than EvoMaster (130). The results clearly show that LlamaRestTest outperforms all other tools, detecting more faults across diverse services.
The most notable detected errors were in the Ohsome and Spotify services. For Ohsome, LlamaRestTest was the only tool to detect faults, identifying 36 errors where all other tools failed. Similarly, LlamaRestTest was the only tool to find faults in Spotify, detecting 8 errors where all others found none. Both if these instances involve processing of complex scenarios. For example, the get /playlists/playlist_id/tracks operation in Spotify’s API requires specific knowledge of how Spotify generates playlist_id. Spotify generates IDs for playlists that are typically 22 characters long, with constraints on the permitted characters and patterns. While many tools fail to generate valid IDs, LlamaREST-EX, due to its fine-tuning, accurately retrieved and generated valid Spotify playlist IDs. As a result, LlamaRestTest was able to successfully execute the get /playlists/playlist_id/tracks operation, triggering a chain of interactions across the service. For instance, after retrieving the International Standard Recording Code (ISRC) from the playlist’s tracks, the mutation process randomly selects an ISRC to use as user_id in the subsequent get /users/user_id/playlists operation. This sequence exposes a hidden dependency conflict, which causes an Internal Server Error.
LlamaRestTest outperforms state-of-the-art REST API testing tools in fault detection by triggering 204 Internal Server Error (500) responses over 10 runs, thus identifying 44 to 74 more errors than the other tools, which illustrates its superior ability to detect internal server errors.
@@ 4.6. RQ5: Ablation Study
Table 7. Ablation study on the impact of key components of LlamaRestTest on coverage achieved.
@@
1. Without Parameter Description 49.9% (-5.9%) 26.9% (-1.4%) 50.8% (-4.5%)
2. Without Server Response 49.7% (-6.1%) 23.9% (-4.4%) 49.0% (-6.3%)
3. Without LlamaREST-IPD 46.6% (-9.2%) 22.1% (-6.2%) 46.9% (-8.4%)
4. Without LlamaREST-EX 52.1% (-3.7%) 23.8% (-4.5%) 52.3% (-3.0%)
To assess the contribution of the key components of LlamaRestTest, we conducted an ablation study by systematically removing specific features and measuring the resulting effect on method, branch, and line coverage across three runs. In particular, we disabled the following features individually to study their effects on coverage: analysis of parameter descriptions, analysis of server response messages, LlamaREST-IPD, and LlamaREST-EX. Table 7 presents the code coverage achieved by LlamaRestTest with and without these components, thus showing how much each feature contributes to its overall performance.
The full LlamaRestTest system achieved method coverage of 55.8%, branch coverage of 28.3%, and line coverage of 55.3%. Removing the parameter description feature resulted in a noticeable drop, with method coverage decreasing by 5.9 percentage points, branch coverage by 1.4 percentage points, and line coverage by 4.5 percentage points. This suggests that including parameter descriptions improves test case generation by providing relevant information for generation of valid inputs. Interestingly, excluding the server response feature caused an even more significant drop, particularly in method coverage, which fell by 6.1 percentage points, followed by branch and line coverage, which decreased by 4.4 and 6.3 percentage points, respectively. While previous studies in this area have focused primarily on enhancing specifications (Kim et al., 2023, 2024b), these results highlight the critical role of server feedback in guiding test case generation and increasing coverage via analysis of dynamic behaviors at runtime.
The removal of LlamaREST-IPD had the most substantial impact on performance, reducing method coverage by 9.2 percentage points, branch coverage by 6.2 percentage points, and line coverage by 8.4 percentage points. This suggests that modeling inter-parameter dependencies is crucial for generating more effective test cases that cover complex interactions between parameters. Finally, excluding LlamaREST-EX resulted in moderate reductions in coverage, with method, branch, and line coverage decreasing by 3.7, 4.5, and 3.0 percentage points, respectively. But LlamaREST-EX still contributes meaningfully to overall test effectiveness, particularly in terms of branch coverage.
The ablation on key components of LlamaRestTest shows that, while each component contributes to the overall performance of LlamaRestTest, server response analysis and inter-parameter dependency modeling are the more critical components, with their removal causing larger drops in code coverage than the other components.